OpenStudy (austinlr12):

A perpendicular bisector,CD , is drawn through point C on AB . If the coordinates of point A are (-3, 2) and the coordinates of point B are (7, 6), the x-intercept of CD is . Point lies on Cd.

OpenStudy (austinlr12):

@Michele_Laino

OpenStudy (adonine):

wht steps do u think we need to do first

OpenStudy (austinlr12):

I have no clue tbh I am in a time rush...

OpenStudy (adonine):

brb

OpenStudy (adonine):

@Michele_Laino

OpenStudy (adonine):

@Atsie

OpenStudy (austinlr12):

@Bigbosssaint21

Atsie (atsie):

@jigglypuff314 I'm really bad at math and don't want to be misguiding, so Jig's can you help out at all? Thanks in advance!! <3

OpenStudy (austinlr12):

Thx Everyone!

OpenStudy (lhgraythorn):

Geometry, eh, just finished the course, let me see if I can help.

jigglypuff314 (jigglypuff314):

This is... complicated - I haven't done geometry in years, so forgive me if I get something wrong your first step should be to find point C that would be the midpoint of AB given A is (-3,2) and B is (7,6) do you know how to find the midpoint? :)

jigglypuff314 (jigglypuff314):



OpenStudy (austinlr12):

And that would be the x-intercept of CD is?

jigglypuff314 (jigglypuff314):

We first have to find where C is in order to find the x-intercept of a line containing C

OpenStudy (austinlr12):

ooh

jigglypuff314 (jigglypuff314):

Do you know how to find the midpoint between the points (-3, 2) and (7, 6) ?

OpenStudy (austinlr12):

im in a time crunch tbh ...

jigglypuff314 (jigglypuff314):

simplify \[(\frac{ 6-2 }{ 2 },\frac{ 7+3 }{ 2 })\]

jigglypuff314 (jigglypuff314):

what do you get?

OpenStudy (austinlr12):

y=x^2 -7x+10

OpenStudy (austinlr12):

the x-intercept of is . Point lies on . NextReset

jigglypuff314 (jigglypuff314):

you follow my steps, we get you to your answer

OpenStudy (austinlr12):

3,0?

jigglypuff314 (jigglypuff314):

try again simplify \[(\frac{ 6-2 }{ 2 },\frac{ 7+3 }{ 2 })\]

OpenStudy (austinlr12):

9,0

jigglypuff314 (jigglypuff314):

let's take this slower then. what is 6 - 2 = ?

OpenStudy (austinlr12):

4

jigglypuff314 (jigglypuff314):

good now divide that by 2

OpenStudy (austinlr12):

2

jigglypuff314 (jigglypuff314):

perfect now what is 7 + 3 = ?

OpenStudy (austinlr12):

10

jigglypuff314 (jigglypuff314):

good! now divide that by 2

OpenStudy (austinlr12):

5

jigglypuff314 (jigglypuff314):

great! so now we have point C is located at (2, 5) then we try to find the equation of a line given the point (2,5) and a slope of -5/2

OpenStudy (austinlr12):

thats not one of my chooses.

jigglypuff314 (jigglypuff314):

I know if we find the equation of that line, then plug in 0 for y and solve for x that gets us our answer so we must find the equation of the line

OpenStudy (austinlr12):

o

jigglypuff314 (jigglypuff314):

5 = (-5/2)(2) + b can you tell me what b is?

OpenStudy (austinlr12):

10

jigglypuff314 (jigglypuff314):

perfect so we have y = (-5/2)x + 10 x-intercept is when y=0 so 0 = (-5/2)x + 10 solve for x and that's your answer

OpenStudy (austinlr12):

4

jigglypuff314 (jigglypuff314):

tah dahhh

OpenStudy (austinlr12):

so what do I do next? because thats not one of my answer chooses

jigglypuff314 (jigglypuff314):

really? what are your answer choices?